Complete the following reations :

A. H2O + H2O

B. HCI + H2O

C. NAOH + H2O

D. NH40H + H2O

E. H2SO4 + H2O

F. CH3COOH + H2O

1. If the pH of NaOH is 12.7, what is its concentration?

2. If the pOH of HCI is 12.7, what is its concentration?
Answer #1

A. H2O + H2O = 2H2O or H3O+ + OH-

B. HCI + H2O = H3O+ + Cl-

C. NAOH + H2O = Na+(aq) + OH-(aq) + H2O(l)

D. NH4OH + H2O = NH4+ + OH- + H2O(l)

E. H2SO4 + H2O = H3O+(aq) + HSO4-(aq)

F. CH3COOH + H2O = CH3COO-(aq) + H3O+(aq)

1. If the pH of NaOH is 12.7, what is its concentration?

pH = 12.7

pOH = 14-12.7 = 1.3

[OH-] = 10^-pOH = 10^-1.3 = 0.0501187

since ratio is 1:1 for NaOH: OH-

then

[NaOH] = 0.0501187 M

2. If the pOH of HCI is 12.7, what is its concentration?

pH = 14-12.7 = 1.3

then

[H+] = 10^-pH = 10^-1.3= 0.0501187 M

since ratio is 1 mol of H+ per 1 mol of HCl

then

[HCl] = 0.0501187 M